OpenStudy (anonymous):

simplify 2x - 24 over 8 answer: 2x - 3

OpenStudy (anonymous):

did you divide it or ?

OpenStudy (anonymous):

if you simplified it i think that answer should be x-12/4

OpenStudy (anonymous):

yes, divide

OpenStudy (anonymous):

x - 12.... that's all i know

OpenStudy (anonymous):

x-12 over 4

OpenStudy (anonymous):

should i simplify?

OpenStudy (anonymous):

2x - 24)/8 = 2(x-12)/8 = (x-12)/4

OpenStudy (anonymous):

\[\frac{2x - 24}{8}=\frac{2(x-12)}{8}=\frac{x-12}{4}\]

OpenStudy (anonymous):

thank you everyone.